how do set up the proportion problem the numbers that are given are 40 inch hips the maximum ideal waist size is 32 inches. What is the maximum ideal waist size for a woman with a 35 inch waist

Answer by stanbon(75887)

how do set up the proportion problem the numbers that are given are 40 inch hips the maximum ideal waist size is 32 inches. What is the maximum ideal waist size for a woman with a 35 inch waist
------
Proportion:
ideal/hip = ideal/hip
---
x/35 = 32/40
---
x = 35(32/40)
x = (7/8)32
x = 7*4
x = 28 inches
======================
Cheers,
Stan H.